OpenStudy (anonymous):

__1___ - ___2__ = ___2__ y2 - 16 y + 4 y - 4 I keep getting the wrong answer while trying to use the LCD (y + 4)(y - 4)

OpenStudy (anonymous):

1/(y^2 -16) = 2/(y+4) +2/(y-4) 1/(y^2 -16) = 2(y-4)+2(y+4) / (y+4)(y-4) 1/(y^2 -16) = 2(y-4)+2(y+4) / (y^2- 16) 1=2(y-4)+2(y+4) 1=2y-8+2y+8 1=4y y=1/4